# MID: A Comprehensive Shore-Based Dataset for Multi-Scale Dense Ship Occlusion and Interaction Scenarios

### v2

@@ Abstract
This paper introduces a maritime ship navigation behavior dataset, referred to as MID-Ship, which is annotated using Oriented Bounding Boxes (OBB) to address the challenges of ship target detection in complex real-world maritime environments. MID-Ship currently contains 5,673 images with a total of 135,884 finely annotated target instances, providing a robust foundation for both supervised and semi-supervised learning tasks. The dataset encompasses a variety of maritime scenarios, including ship encounters in diverse weather conditions, docking maneuvers, small target aggregation, and partial occlusions. These scenarios contribute to filling the gap in existing datasets (e.g., HRSID, SSDD, NWPU-10) by addressing the lack of data on complex situations such as occlusion and dense interactions, which are particularly relevant in the context of advanced technologies such as autonomous ships. All images in MID-Ship are sourced from high-definition video clips of real-world maritime navigation across 43 distinct navigable water areas. The dataset is further enriched through the inclusion of multiple weather and lighting conditions (e.g., rainy, cloudy, foggy days) and manually supplemented annotations, enhancing its diversity and ensuring that models trained on MID-Ship are better equipped to handle varying natural environments. This diversity also ensures its applicability to the real-world demands of busy ports and dense maritime regions. Using MID-Ship, we conducted a comprehensive evaluation of 10 detection algorithms, presenting a detailed analysis and comparison of their performance. This study includes: 1) an in-depth analysis of the dataset, 2) a showcase of detection results from various models, and 3) a comparison of baseline algorithms, with particular focus on their ability to detect dense occlusions. The results suggest that MID-Ship holds significant potential for advancing research in intelligent maritime traffic monitoring systems and autonomous vessel navigation, with important implications for future applications in maritime environment and safety situational awareness. Our dataset will be released at https://github.com/VirtualNew/MID_DataSet.
@@ I Introduction
In the context of the global development of smart ships and maritime port traffic management, the accuracy of target detection, tracking, and trajectory prediction for ships in complex scenarios is crucial for ensuring port safety, enhancing logistics efficiency, and reducing maritime accidents [1]. With the rapid growth of the global shipping industry, ports have increasingly become hubs of international trade [2]. Therefore, ensuring the safe navigation and efficient operation of ships entering and leaving ports in complex environments is especially important. Modern ports are typically equipped with numerous surveillance cameras to monitor the entry, exit, and activities of ships in real time [3]. However, the complexity of port environments characterized by a high number of ships, overlapping navigation paths, and diverse ship sizes and types—poses significant challenges for ship detection and tracking [4]. In densely populated areas, smaller ships are particularly prone to being obscured or confused with other ships, further complicating detection and tracking tasks [5]. Addressing these issues is not only vital for the operational efficiency of ports but also plays a key role in reducing the incidence of maritime accidents.
@@
Based on the aforementioned issues and the advantages of optical images, this paper presents the MID-Ship, constructed from visible light video images and demonstrating significant benefits. The dataset was meticulously annotated and processed over three months by four professionals in the port sector, ensuring high accuracy and detail. This high-quality annotation provides a reliable foundation for supervised learning, significantly enhancing model performance. Moreover, the detailed labeling system offers substantial potential for semi-supervised learning tasks, effectively guiding models to learn from a limited amount of labeled data and increasing the utilization of unlabeled data. Consequently, the MID-Ship is not only suitable for supervised learning tasks but also serves as an ideal experimental platform for semi-supervised learning research. It provides crucial data support for fields such as small target occlusion detection and ship tracking in maritime traffic management.
The significance of this research lies in the construction of the MID-Ship, which fills the gap in existing datasets regarding multi-scene dense occlusion targets, particularly providing crucial support for small target detection. This dataset encompasses dense ship scenarios and complex interaction behaviors—such as overtaking, crossing, and meeting—ensuring the precise recognition of small ships by detection algorithms in challenging environments, thereby offering robust technical assurance for the safe management of ports and maritime traffic. Additionally, the MID-Ship serves as an ideal testing platform for multi-object tracking tasks, with detailed annotations for scenes with severe occlusion or complex ship behaviors, significantly reducing the issue of target loss during tracking. Furthermore, the dataset provides essential support for the development of trajectory prediction models, covering a variety of ship interaction scenarios and possessing immense application value, especially in preventing ship collisions and managing complex behaviors. The construction of the MID-Ship offers critical support for optimizing detection, tracking, and trajectory prediction technologies, aiding in improving the performance of these visual techniques in complex port environments. By encompassing intricate ship interactions and dense occlusion scenarios, this dataset becomes an important resource for the study of unmanned ship technology, supporting the development and validation of autonomous navigation, obstacle avoidance, and trajectory prediction algorithms in complex marine environments, thus providing a vital data foundation for the practical application of unmanned maritime traffic systems.
@@ III Data Collection
We constructed the MID-Ship based on images captured by video monitoring systems deployed in real environments. Fig. 1 shows the video capture device used, with specific parameters listed in Table I. Three types of high-definition cameras were installed near open water areas of the port and in densely navigated narrow channels, covering a total area of 41 square kilometers. These cameras provide high-quality monitoring videos (1920×1080 pixels) from which the images for our proposed dataset are extracted. Specifically, one high-definition electro-optical device was installed on a platform at a height of 100 meters above the open water area, capturing images of ships entering and leaving the port, including a visible light zoom camera and a controllable rotating gimbal. In the densely navigated narrow channel area, one variable focal length camera and one panoramic fisheye camera were installed on a building at a height of 118 meters to capture images of ships engaged in crossing, overtaking, and encountering maneuvers.
@@ III-A Visible Light Video Image Production
In addition, we systematically organized the images and their annotation files in MID-Ship, dividing them into training, validation, and testing sets. We also included conversion scripts for various annotation formats in the toolkit to support the needs of different detection networks. The specific file structure of MID-Ship is detailed in Fig. 5.
Figure 5: Overall file structure of MID-Ship.
@@ III-B4 The diversity of shapes and tonnages of ships navigating on the sea
Figure 6: Diversity of MID-Ship.
Figure 7: Examples images of different weather variations in MID-Ship. The 8 sub-images correspond to 8 different weather scenes.
Figure 8: Example images of water droplets appearing on cameras in MID-Ship.
Figure 9: Scatter plot of instance width vs height in MID-Ship.
Figure 10: Proportion chart of different scale instances in MID-Ship.
Figure 11: The example images in the first to second rows show six different sub viewpoints of the same ship in MID-Ship in a panoramic view. The third and fourth rows are example images of scale changes in two scenes, with close-up views, respectively. The two rows of images show scale changes in two different scenes.
Figure 12: Distribution of ship instances of each image in MID-Ship.
IV DESIGN AND ANALYSIS OF MID-Ship
To highlight the advantages of MID-Ship, we conduct an in-depth analysis from six key aspects: weather conditions, target scale, target aspect ratio, background interference, target occlusion, and ship collision avoidance. These factors significantly impact the performance of object detection, reflecting the diversity and application potential of the dataset.
@@ IV-A Weather Variation
Due to the diversity of weather conditions, which provides richer visual effects, the image dataset becomes more representative and can better reflect various real-world scenes. Therefore, we analyzed the types of weather conditions in MID-Ship and categorized and analyzed the data for each type of scene. We classified weather conditions into eight categories: silty, waves, rainy, fog, overexposed, dark, fuzzy, and color-distorted, as shown in Fig. 7.
@@ IV-B Scale Variation
Scale variation is an important characteristic, reflecting how ships of different sizes appear in images. Due to varying shooting conditions and perspectives, the same vessel may present different scales across images. This variation can stem from factors such as shooting distance, lens focal length, and environmental conditions. To explore scale variation in MID-Ship, we first examined the overall scale distribution in the dataset, as shown in Table II. Fig. 9 and 10 illustrate the scale of labeled targets more vividly through scatter plots and proportion charts. We then broadly categorized the shooting scenes into two types: panoramic and close-up. Fig. 11 describes the scale variation in both close-up and distant views.
Fig. 12 shows the distribution of ship instances per image in MID-Ship, where most images contain between 5-25 instances, and a few contain 35-55 instances. On average, each image includes 23.953 ship instances, with a maximum of 81 and a minimum of 1 instance per image.
Figure 13: AR distribution in MID-Ship.
TABLE II: Scale variation statistics in MID-Ship.
@@ IV-C AR of Instance Variations
The aspect ratio (AR) of ship objects is an important feature, typically ranging from 3 to 6, and is one of the key characteristics for ship classification and identification tasks. For deep learning-based methods, AR serves as a crucial parameter when designing anchor-based models, such as Faster R-CNN [29]. Fig. 13 illustrates the distribution of ship aspect ratios in MID-Ship.
@@ IV-D Viewpoint Variation
Figure 15: Five main different backgrounds in MID-Ship.
Figure 16: Sixteen different backgrounds in MID-Ship based on the number of the videos.
@@ IV-E Background Variation
The background diversity can be investigated based on the cameras deployed at different locations. To this end, we conducted an analysis of MID-Ship based on shooting angles and scenes, categorizing the backgrounds into five main types, as shown in Fig. 15. Additionally, based on the number of video files captured, we classified the backgrounds into 16 categories and counted the number of images under each background, as shown in Fig. 16.
@@ IV-F Occlusion Variation
To investigate the types of occlusion in MID-Ship, we grouped all images with occlusion across different scenes and analyzed the degree of occlusion, as shown in Table III. Our findings reveal that 16% of the dataset contains varying levels of occlusion, with some images showing over 90% occlusion or even complete occlusion. This makes MID-Ship a particularly challenging dataset for handling occlusions. Fig. 17 provides a visual representation of the distribution of ships under different levels of occlusion.
@@
      Method       P       R       mAP.50(%)       mAP.50:.95(%)       Param(m)       Flops(G)
@@ IV-G Collision Variation
To explore the types of collision avoidance scenarios in MID-Ship, we grouped all images with collision avoidance situations from different scenes and analyzed them. We classified the collision avoidance scenarios into three main types: crossing, overtaking, and head-on encounter. Crossing refers to two ships’ paths forming a relative intersection, usually sailing on different courses that cross each other. Overtaking refers to one ship approaching and passing another from behind. Head-on encounter refers to two ships sailing directly toward each other, facing head-on. These scenarios are illustrated in Fig. 18.

### v3

@@ Abstract
This paper introduces a maritime ship navigation behavior dataset, referred to as MID, which is annotated using Oriented Bounding Boxes (OBB) to address the challenges of ship target detection in complex real-world maritime environments. MID currently contains 5,673 images with a total of 135,884 finely annotated target instances, providing a robust foundation for both supervised and semi-supervised learning tasks. The dataset encompasses a variety of maritime scenarios, including ship encounters in diverse weather conditions, docking maneuvers, small target aggregation, and partial occlusions. These scenarios contribute to filling the gap in existing datasets (e.g., HRSID, SSDD, NWPU-10) by addressing the lack of data on complex situations such as occlusion and dense interactions, which are particularly relevant in the context of advanced technologies such as autonomous ships. All images in MID are sourced from high-definition video clips of real-world maritime navigation across 43 distinct navigable water areas. The dataset is further enriched through the inclusion of multiple weather and lighting conditions (e.g., rainy, cloudy, foggy days) and manually supplemented annotations, enhancing its diversity and ensuring that models trained on MID are better equipped to handle varying natural environments. This diversity also ensures its applicability to the real-world demands of busy ports and dense maritime regions. Using MID, we conducted a comprehensive evaluation of 10 detection algorithms, presenting a detailed analysis and comparison of their performance. This study includes: 1) an in-depth analysis of the dataset, 2) a showcase of detection results from various models, and 3) a comparison of baseline algorithms, with particular focus on their ability to detect dense occlusions. The results suggest that MID holds significant potential for advancing research in intelligent maritime traffic monitoring systems and autonomous vessel navigation, with important implications for future applications in maritime environment and safety situational awareness. Our dataset will be released at https://github.com/VirtualNew/MID_DataSet.
@@ I Introduction
In the context of the global development of smart ships and maritime port traffic management, the accuracy of target detection, tracking, and trajectory prediction for ships in complex scenarios is crucial for ensuring port safety, enhancing logistics efficiency, and reducing maritime accidents [1]. With the rapid growth of the global shipping industry, ports have increasingly become hubs of international trade [2]. Therefore, ensuring the safe navigation and efficient operation of ships entering and leaving ports in complex environments is especially important. Modern ports are typically equipped with numerous surveillance cameras to monitor the entry, exit, and activities of ships in real time [3]. However, the complexity of port environments characterized by a high number of ships, overlapping navigation paths, and diverse ship sizes and types poses significant challenges for ship detection and tracking [4]. In densely populated areas, smaller ships are particularly prone to being obscured or confused with other ships, further complicating detection and tracking tasks [5]. Addressing these issues is not only vital for the operational efficiency of ports but also plays a key role in reducing the incidence of maritime accidents.
@@
Based on the aforementioned issues and the advantages of optical images, this paper presents the MID, constructed from visible light video images and demonstrating significant benefits. The dataset was meticulously annotated and processed over three months by four professionals in the port sector, ensuring high accuracy and detail. This high-quality annotation provides a reliable foundation for supervised learning, significantly enhancing model performance. Moreover, the detailed labeling system offers substantial potential for semi-supervised learning tasks, effectively guiding models to learn from a limited amount of labeled data and increasing the utilization of unlabeled data. Consequently, the MID is not only suitable for supervised learning tasks but also serves as an ideal experimental platform for semi-supervised learning research. It provides crucial data support for fields such as small target occlusion detection and ship tracking in maritime traffic management.
The significance of this research lies in the construction of the MID, which fills the gap in existing datasets regarding multi-scene dense occlusion targets, particularly providing crucial support for small target detection. This dataset encompasses dense ship scenarios and complex interaction behaviors—such as overtaking, crossing, and meeting—ensuring the precise recognition of small ships by detection algorithms in challenging environments, thereby offering robust technical assurance for the safe management of ports and maritime traffic. Additionally, the MID serves as an ideal testing platform for multi-object tracking tasks, with detailed annotations for scenes with severe occlusion or complex ship behaviors, significantly reducing the issue of target loss during tracking. Furthermore, the dataset provides essential support for the development of trajectory prediction models, covering a variety of ship interaction scenarios and possessing immense application value, especially in preventing ship collisions and managing complex behaviors. The construction of the MID offers critical support for optimizing detection, tracking, and trajectory prediction technologies, aiding in improving the performance of these visual techniques in complex port environments. By encompassing intricate ship interactions and dense occlusion scenarios, this dataset becomes an important resource for the study of unmanned ship technology, supporting the development and validation of autonomous navigation, obstacle avoidance, and trajectory prediction algorithms in complex marine environments, thus providing a vital data foundation for the practical application of unmanned maritime traffic systems.
@@ III Data Collection
We constructed the MID based on images captured by video monitoring systems deployed in real environments. Fig. 1 shows the video capture device used, with specific parameters listed in Table I. Three types of high-definition cameras were installed near open water areas of the port and in densely navigated narrow channels, covering a total area of 41 square kilometers. These cameras provide high-quality monitoring videos (1920×1080 pixels) from which the images for our proposed dataset are extracted. Specifically, one high-definition electro-optical device was installed on a platform at a height of 100 meters above the open water area, capturing images of ships entering and leaving the port, including a visible light zoom camera and a controllable rotating gimbal. In the densely navigated narrow channel area, one variable focal length camera and one panoramic fisheye camera were installed on a building at a height of 118 meters to capture images of ships engaged in crossing, overtaking, and encountering maneuvers.
@@ III-A Visible Light Video Image Production
In addition, we systematically organized the images and their annotation files in MID, dividing them into training, validation, and testing sets. We also included conversion scripts for various annotation formats in the toolkit to support the needs of different detection networks. The specific file structure of MID is detailed in Fig. 5.
Figure 5: Overall file structure of MID.
@@ III-B4 The diversity of shapes and tonnages of ships navigating on the sea
Figure 6: Diversity of MID.
Figure 7: Examples images of different weather variations in MID. The 8 sub-images correspond to 8 different weather scenes.
Figure 8: Example images of water droplets appearing on cameras in MID.
Figure 9: Scatter plot of instance width vs height in MID.
Figure 10: Proportion chart of different scale instances in MID.
Figure 11: The example images in the first to second rows show six different sub viewpoints of the same ship in MID in a panoramic view. The third and fourth rows are example images of scale changes in two scenes, with close-up views, respectively. The two rows of images show scale changes in two different scenes.
Figure 12: Distribution of ship instances of each image in MID.
IV DESIGN AND ANALYSIS OF MID
To highlight the advantages of MID, we conduct an in-depth analysis from six key aspects: weather conditions, target scale, target aspect ratio, background interference, target occlusion, and ship collision avoidance. These factors significantly impact the performance of object detection, reflecting the diversity and application potential of the dataset.
@@ IV-A Weather Variation
Due to the diversity of weather conditions, which provides richer visual effects, the image dataset becomes more representative and can better reflect various real-world scenes. Therefore, we analyzed the types of weather conditions in MID and categorized and analyzed the data for each type of scene. We classified weather conditions into eight categories: silty, waves, rainy, fog, overexposed, dark, fuzzy, and color-distorted, as shown in Fig. 7.
@@ IV-B Scale Variation
Scale variation is an important characteristic, reflecting how ships of different sizes appear in images. Due to varying shooting conditions and perspectives, the same vessel may present different scales across images. This variation can stem from factors such as shooting distance, lens focal length, and environmental conditions. To explore scale variation in MID, we first examined the overall scale distribution in the dataset, as shown in Table II. Fig. 9 and 10 illustrate the scale of labeled targets more vividly through scatter plots and proportion charts. We then broadly categorized the shooting scenes into two types: panoramic and close-up. Fig. 11 describes the scale variation in both close-up and distant views.
Fig. 12 shows the distribution of ship instances per image in MID, where most images contain between 5-25 instances, and a few contain 35-55 instances. On average, each image includes 23.953 ship instances, with a maximum of 81 and a minimum of 1 instance per image.
Figure 13: AR distribution in MID.
TABLE II: Scale variation statistics in MID.
@@ IV-C AR of Instance Variations
The aspect ratio (AR) of ship objects is an important feature, typically ranging from 3 to 6, and is one of the key characteristics for ship classification and identification tasks. For deep learning-based methods, AR serves as a crucial parameter when designing anchor-based models, such as Faster R-CNN [29]. Fig. 13 illustrates the distribution of ship aspect ratios in MID.
@@ IV-D Viewpoint Variation
Figure 15: Five main different backgrounds in MID.
Figure 16: Sixteen different backgrounds in MID based on the number of the videos.
@@ IV-E Background Variation
The background diversity can be investigated based on the cameras deployed at different locations. To this end, we conducted an analysis of MID based on shooting angles and scenes, categorizing the backgrounds into five main types, as shown in Fig. 15. Additionally, based on the number of video files captured, we classified the backgrounds into 16 categories and counted the number of images under each background, as shown in Fig. 16.
@@ IV-F Occlusion Variation
To investigate the types of occlusion in MID, we grouped all images with occlusion across different scenes and analyzed the degree of occlusion, as shown in Table III. Our findings reveal that 16% of the dataset contains varying levels of occlusion, with some images showing over 90% occlusion or even complete occlusion. This makes MID a particularly challenging dataset for handling occlusions. Fig. 17 provides a visual representation of the distribution of ships under different levels of occlusion.
@@
      Method       P       R       $\mathrm{mAP}_{.50}$(%)       $\mathrm{mAP}_{.50:.95}$(%)       Param(m)       Flops(G)
@@ IV-G Collision Variation
To explore the types of collision avoidance scenarios in MID, we grouped all images with collision avoidance situations from different scenes and analyzed them. We classified the collision avoidance scenarios into three main types: crossing, overtaking, and head-on encounter. Crossing refers to two ships’ paths forming a relative intersection, usually sailing on different courses that cross each other. Overtaking refers to one ship approaching and passing another from behind. Head-on encounter refers to two ships sailing directly toward each other, facing head-on. These scenarios are illustrated in Fig. 18.